# 3D VR Sketch Guided 3D Shape Prototyping and Exploration

### v1

@@ Abstract
3D shape modeling is labor-intensive and time-consuming and requires years of expertise. Recently, 2D sketches and text inputs were considered as conditional modalities to 3D shape generation networks to facilitate 3D shape modeling. However, text does not contain enough fine-grained information and is more suitable to describe a category or appearance rather than geometry, while 2D sketches are ambiguous, and depicting complex 3D shapes in 2D again requires extensive practice. Instead, we explore virtual reality sketches that are drawn directly in 3D. We assume that the sketches are created by novices, without any art training, and aim to reconstruct physically-plausible 3D shapes. Since such sketches are potentially ambiguous, we tackle the problem of the generation of multiple 3D shapes that follow the input sketch structure. Limited in the size of the training data, we carefully design our method, training the model step-by-step and leveraging multi-modal 3D shape representation. To guarantee the plausibility of generated 3D shapes we leverage the normalizing flow that models the distribution of the latent space of 3D shapes. To encourage the fidelity of the generated 3D models to an input sketch, we propose a dedicated loss that we deploy at different stages of the training process. We plan to make our code publicly available.
Figure 1: Given a VR (Virtual Reality) sketch input, we generate 3D shape samples that satisfy three requirements: (1 - fidelity) reconstructed shapes follow the overall structure of a quick VR sketch; (2 - diversity) reconstructed shapes contain some diversity in shape details: such as a hollow or solid backrest, and (3 - plausibility) reconstructions favor physically-plausible 3D shapes.
@@ 1 Introduction
The demand for convenient tools for 3D content creation constantly grows as the creation of virtual worlds becomes an integral part of various fields such as architecture and cinematography. Recently, several works have demonstrated how text and image priors can be used to create 3D models [44, 29, 28, 16, 27]. However, it is universally accepted that text is much less expressive or precise than a 2D freehand sketch in conveying spatial or geometric information [54, 12, 45]. Therefore, many works focus on sketch-based modeling from 2D sketches [39, 5, 4, 6, 32, 26, 13, 47, 59, 58, 56, 18, 22] as a convenient tool for creating virtual 3D content. Yet, 2D sketches are ambiguous, and depicting a complex 3D shape in 2D requires substantial sketching expertise. As Virtual Reality (VR) headsets and associated technologies progress [21, 20, 36], more and more works consider 3D VR sketch as an input modality in the context of 3D modeling [53, 55, 43] and retrieval [24, 25, 23, 52, 34, 35, 33]. 3D VR sketches are drawn directly in 3D and therefore contain less ambiguity.
Existing works on 3D shape modeling assume rather carefully created inputs and focus more on the interface and logistics of the sketching process. In this paper, we propose a method for 3D shape modeling using 3D VR sketching, but unlike existing works, we aim at rapidly drawn sketches that can be created by anyone and can potentially be quite inaccurate. Such sketches are also more appropriate for the earlier 3D shape idea exploration stage – commonly referred to as the ideation stage. Assuming that the input VR sketch is sparse and potentially ambiguous, we argue that it is not enough to return one 3D shape model, but rather, we want to ensure that multiple shape variations can be generated resembling the input sketch, as shown in Fig. 1. The user can then either directly pick one of the models, or refine the design given the visualized 3D shapes, or multiple shapes can be used in some physical simulation process to select an optimal shape within the constraints provided in the form of a VR sketch.
@@
One of the challenges of working with freehand VR sketches is the lack of datasets, the limited size of such datasets, and the nature of such datasets. We tackle the associated challenge by carefully designing our method and training it step-by-step to obtain robust performance. Sketch-based modeling requires datasets with instance-level correspondence between 3D models and sketches. We are aware of only one fine-grained dataset of VR sketches by Luo et al. [35], which we use in this work. The additional challenge of working with this data comes from the misalignments between sketches and 3D shapes. Luo et al. [35] let participants sketch in an area different from the one where the reference 3D shape is displayed. This allows them to model the scenario of sketching from memory or imagination, however, results in a lack of alignment between 3D shapes and sketches, as shown in Fig. 2. Considering the misalignment of sketches and shapes in the dataset, and the ambiguity of the VR sketches, we aim to generate shapes with good fidelity to an input sketch, rather than the reference shape.
@@
  • We, for the first time, study the problem of conditional 3D shape generation from rapid and sparse 3D VR sketches, and carefully design our method to tackle the problem of (1) limited data, (2) misalignments between sketches and 3D shapes and (3) abstract nature of freehand sketches.
@@ 2.1 Shape reconstruction and retrieval from 3D sketches
2.2 3D models generation
In addition to maximizing the fidelity to the input sketch, we also aim to generate a set of shapes that are distinctive from each other. This diversity allows users to efficiently explore the design space of shapes resembling their initial sketch.
@@ 2.2.2 Conditional shape generation
Sketch-based and image-based reconstruction methods typically focus on the generation of only one output result for each input, while we aim at the generation of multiple 3D models. Meanwhile, in the point cloud completion task, it is typically to infer the missing parts from the observed parts and generate multiple possible completion results. Their task, however, differs from our goal because we do not want the network to synthesize non-existent parts, but only to create various shapes that match the sparse freehand sketch taking into account how humans might abstract 3D shapes. This is also the reason why the autoregressive approaches, such as [46, 37, 51], are not suitable for our problem.
Text-guided 3D shape generation shares similar ambiguity property as VR sketch-guided, i.e., diverse results may match the same input text. CLIP-Forge [44] employs pre-trained visual-textual embedding model CLIP to bridge text and 3D domains, and uses conditional normalizing flow to model the conditional distribution of latent shape representation given text or image embeddings. Zhengzhe et al. [30] introduce shape IMLE (Implicit Maximum Likelihood Estimation) to boost results diversity while utilizing a cyclic loss to encourage consistency. In our work, we condition on a VR sketch rather than text, and aim to obtain diverse 3D shapes that follow the input sketch structure.
@@ 3 Method
We present a conditional generation method that generates physically-plausible shapes conditioned on abstract, sparse, and inaccurate free-hand VR sketches. Our goal is to enforce the generation to stay close to the input sketch (sketch fidelity) while providing sufficient diversity of 3D reconstructions.
@@ 3.1 Shape decoder
We represent 3D shapes using signed distance functions (SDFs), as one of the most common 3D shape representations. This representation is limited to watertight shapes: to support non-watertight surfaces, one can switch to, for example, a recently proposed tri-plane 3D shape representation. But without loss of generality, here we can assume that our meshes are watertight.
A SDF is a continuous function of the form:
@@
The auto-decoder is trained by minimizing an $L_{1}$ loss between the ground truth and predicted signed distance values. The decoder
$\mathrm{D}_{\theta}([\boldsymbol{e}_{f},p_{i}])=s,\ p_{i}\in\mathbb{R}^{3},\ s\in\mathbb{R}$ (2)
takes as input 3D shape latent code $\boldsymbol{e}_{f}$ and 3D point coordinates $p_{i}$; [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents a concatenation operation. Once the decoder is trained, the decoder parameters $\theta$ are frozen in our work.
@@
$\boldsymbol{\hat{e}}_{f}=\underset{\boldsymbol{e}_{f}}{\arg\min}\sum_{\left(\boldsymbol{x}_{j},\boldsymbol{s}_{j}\right)\in X}\mathcal{L}\left(f_{\theta}\left(\boldsymbol{e}_{f},\boldsymbol{x}_{j}\right),s_{j}\right)+\frac{1}{\sigma^{2}}\|\boldsymbol{e}_{f}\|_{2}^{2}$ (3)
where the latent vector $e_{f}$ is initialized randomly. We treat the estimated latent vector $\boldsymbol{\hat{e}}_{f}$ as a ground truth shape embedding and denote it as $\boldsymbol{e}_{f}$ for simplicity.
@@
Due to the sparsity of sketch inputs, we represent them as point clouds. However, if we only use sketches for training, we obtain very poor generalization to the test set. Therefore, we exploit a joint encoder for sketches and 3D shapes, using a PointNet++ [42] encoder as a backbone. The encoder $\mathrm{E}_{\phi}(\cdot):\mathbb{R}^{N_{s}\times 3}\rightarrow\mathbb{R}^{256}$ embeds randomly sampled points from input shape surface $f\in\mathbb{R}^{N_{s}\times 3}$ or sketch strokes $g\in\mathbb{R}^{N_{s}\times 3}$ to a feature vector $\boldsymbol{e}_{f}^{\prime}\in\mathbb{R}^{256}$ and $\boldsymbol{e}^{\prime}_{g}\in\mathbb{R}^{256}$, respectively.
@@
First loss, $\mathcal{L}_{\text{shape}}$, operates only on 3D shape inputs, and ensures that the full network $\mathrm{D}_{\theta}([\mathrm{E}_{\phi},p_{i}])$ functions as a 3D shape autoencoder. First, we minimize the sum of $L_{1}$ losses between the predicted and ground truth SDF values of sampled points $p_{i}\in\mathbb{R}^{3}$:
$\mathcal{L}_{\text{SDF}}(\phi)=\frac{1}{N_{s}}\sum_{p_{i}}|\mathrm{D}_{\theta}([\mathrm{E}_{\phi}(f),p_{i}])-s(p_{i})|$ (5)
where the decoder parameters $\theta$ are from our auto-decoder and are frozen when training the sketch/shape encoder and $s(\cdot)$ denotes ground-truth 3D shape SDF values.
Additionally, we ensure that the 3D shape embedding $\boldsymbol{e}^{\prime}_{f}=\mathrm{E}_{\phi}(f)$ maps directly to a latent representation of a 3D shape $\boldsymbol{e}_{f}$, computed with Eq. 3:
$\mathcal{L}_{f\text{-}L_{1}}(\phi)=|\boldsymbol{e}_{f}-\mathrm{E}_{\phi}(f)|.$ (6)
We also found that adding contrastive latent term loss increases performance. Let’s assume that our mini-batch consists of $N_{f}$ shapes. First, we obtain latent representations $\{e_{f}^{1},\dots,e_{f}^{N_{f}}\}$ of 3D shapes, using Eq. 3. Then, we formulate our contrastive loss term as follows:
$\mathcal{L}_{\text{f-NCE}}(\phi)=-\sum_{i=1}^{N_{f}}\left[\log\frac{\exp\left(-\left|\mathrm{E}_{\phi}(f^{i})-e_{f}^{i}\right|\right)}{\sum_{j=1}^{N_{f}}\exp\left(-\left|\mathrm{E}_{\phi}(f^{i})-e_{f}^{j}\right|\right)}\right],$ (7)
where $f^{i}$ is the point cloud representation of the $i$-th shape.
@@
$\mathcal{L}_{\text{shape}}=\mathcal{L}_{\text{SDF}}(\phi)+\mathcal{L}_{f\text{-}L_{1}}(\phi)+\mathcal{L}_{f\text{-NCE}}(\phi).$ (8)
@@ 3.2.2 Sketch loss
Given the misalignment between sketches and reference 3D shapes in the existing dataset [34], as we show in Fig. 2, we aim for the reconstruction result to stay close to the sketch input. To achieve this goal, we design a sketch loss $\mathcal{L}_{\text{sketch}}$.
Since a sketch is a sparse representation of a 3D shape, we assume that the sketch stroke points $g\in\mathbb{R}^{N_{s}\times 3}$ represent shape surface points. Therefore, the reconstructed SDF values at those points should be close to zero. Formally, we define this loss as follows:
$\mathcal{L}_{\text{sketch}}(\phi)=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}|D(\mathrm{E}_{\phi}(g),g_{i})|,$ (9)
where $g_{i}$ is the $i$-th sample points from the input sketch.
@@ 3.2.3 Sketch-shape latent space alignment
The considered so far losses do not guarantee that there is a meaningful mapping between sketches and 3D shapes’ latent representations. Therefore, we design additional losses encouraging an alignment in the feature space.
@@
$\mathcal{L}_{g\text{-NCE}}(\phi)=-\sum_{i=1}^{N_{g}}\left[\log\frac{\exp\left(-\left|\mathrm{E}_{\phi}(g^{i})-e_{f}^{i}\right|\right)}{\sum_{j=1}^{N_{f}}\exp\left(-\left|\mathrm{E}_{\phi}(g^{i})-e_{f}^{j}\right|\right)}\right],$ (10)
where $N_{g}$ is the number of sketches and $N_{f}$ is the number of shapes in the mini-batch. This loss pulls the encodings of a sketch and a reference shape closer than the encodings of a sketch and non-matching 3D shapes.
Additionally, we minimize the $L_{1}$ distance between the sketch embedding $\boldsymbol{e}^{\prime}_{g}=E_{\phi}(g^{i})$ to the ground-truth shape latent code $e_{f}$:
$\mathcal{L}_{g\text{-}L_{1}}(\phi)=|\boldsymbol{e}_{f}-E_{\phi}(g^{i})|.$ (11)
@@
$\mathcal{L}_{\text{align}}=\mathcal{L}_{g\text{-}L_{1}}+\mathcal{L}_{g\text{-NCE}}$ (12)

### v6

@@ Abstract
3D shape modeling is labor-intensive, time-consuming, and requires years of expertise. To facilitate 3D shape modeling, we propose a 3D shape generation network that takes a 3D VR sketch as a condition. We assume that sketches are created by novices without art training and aim to reconstruct geometrically realistic 3D shapes of a given category. To handle potential sketch ambiguity, our method creates multiple 3D shapes that align with the original sketch’s structure. We carefully design our method, training the model step-by-step and leveraging multi-modal 3D shape representation to support training with limited training data. To guarantee the realism of generated 3D shapes, we leverage the normalizing flow that models the distribution of the latent space of 3D shapes. To encourage the fidelity of the generated 3D shapes to an input sketch, we propose a dedicated loss that we deploy at different stages of the training process. The code is available at https://github.com/Rowl1ng/3Dsketch2shape.
@@ 1 Introduction
The demand for convenient tools for 3D content creation constantly grows as the creation of virtual worlds becomes an integral part of various fields such as architecture and cinematography. Recently, several works have demonstrated how text and image priors can be used to create 3D shapes [46, 30, 29, 16, 28]. However, it is universally accepted that text is much less expressive or precise than a 2D freehand sketch in conveying spatial or geometric information [56, 12, 47]. Therefore, many works focus on sketch-based modeling from 2D sketches [40, 5, 4, 6, 33, 27, 13, 49, 61, 60, 58, 18, 23] as a convenient tool for creating virtual 3D content. Yet, 2D sketches are ambiguous, and depicting a complex 3D shape in 2D requires substantial sketching expertise. As Virtual Reality (VR) headsets and associated technologies progress [22, 21, 37], more and more works consider 3D VR sketch as an input modality in the context of 3D modeling [55, 57, 45] and retrieval [25, 26, 24, 54, 35, 36, 34]. Firstly, 3D VR sketches are drawn directly in 3D and therefore provide a more immersive and intuitive design experience. Secondly, 3D VR sketches offer a natural way to convey volumetric shapes and spatial relationships. Moreover, the use of 3D VR sketches aligns with advancements in virtual reality technology, making the process of sketching and designing more future-proof and adaptable to emerging technologies.
Figure 1: Given a VR (Virtual Reality) sketch input, we generate 3D shape samples that satisfy three requirements: (1 - fidelity) reconstructed shapes follow the overall structure of a quick VR sketch; (2 - diversity) reconstructed shapes contain some diversity in shape details: such as a hollow or solid backrest, and (3 - realism) reconstructions favor geometrically realistic 3D shapes of a given category.
Existing works on 3D shape modeling assume carefully created inputs and focus primarily on the interface and logistics of the sketching process. In this paper, we introduce a novel method for 3D shape modeling that utilizes 3D VR sketching. Our method does not require professional sketch training or detailed sketches and is trained and tested on a dataset of sketches created by participants without art experience. This approach ensures that our model can accommodate a diverse range of users and handle sketches that might be less polished or precise, making it more accessible and practical for real-world applications. Considering the sparsity of VR sketches, a single 3D shape model may not match the user’s intention. Therefore, we advocate for generating multiple shape variations that closely resemble the input sketch, as demonstrated in Fig. 1. The user can then either directly pick one of the models, or refine the design given the visualized 3D shapes, or multiple shapes can be used in some physical simulation process to select the optimal shape within the constraints of the VR sketch.
@@
Working with freehand VR sketches presents several challenges due to the lack of datasets. We are aware of only one fine-grained dataset of VR sketches by Luo et al. [36], which we use in this work. The challenge of working with this data comes from its limited size and the misalignment between sketches and 3D shapes. Luo et al. [36] let participants sketch in an area different from the one where the reference 3D shape is displayed. This allows to model the scenario of sketching from memory or imagination, however, results in a lack of alignment between 3D shapes and sketches, as shown in Fig. 2. Considering the misalignment of sketches and shapes in the dataset, and the ambiguity of the VR sketches, we aim to generate shapes with good fidelity to an input sketch, rather than the reference shape.
@@
  • We, for the first time, study the problem of conditional 3D shape generation from rapid and sparse 3D VR sketches and carefully design our method to tackle the problem of (1) limited data, (2) misalignments between sketches and 3D shapes and (3) abstract nature of freehand sketches.
@@ 2.1 Shape reconstruction and retrieval from 3D sketches
2.2 3D shapes generation
In addition to maximizing the fidelity to the input sketch, we aim to generate a set of shapes that are distinctive from each other. This diversity allows users to efficiently explore the design space of shapes resembling their initial sketch.
@@ 2.2.2 Conditional shape generation
Sketch-/image-based reconstruction methods typically focus on the generation of only one output result for each input, while we aim at the generation of multiple 3D shapes. Meanwhile, in the point cloud completion task, it is typically to infer the missing parts from the observed parts and generate multiple possible completion results. Their task, however, differs from our goal as we do not want the network to synthesize non-existent parts, but only to create various shapes that match the sparse freehand sketch taking into account how humans might abstract 3D shapes. This is also the reason why the autoregressive approaches, such as [48, 38, 53], are not suitable for our problem.
Text-guided 3D shape generation shares similar ambiguity properties as VR sketch-guided, i.e., diverse results may match the same input text. CLIP-Forge [46] employs pre-trained visual-textual embedding model CLIP to bridge text and 3D domains, and uses conditional normalizing flow to model the conditional distribution of latent shape representation given text or image embeddings. Zhengzhe et al. [31] introduce shape IMLE (Implicit Maximum Likelihood Estimation) to boost results diversity while utilizing a cyclic loss to encourage consistency. In our work, we condition on a VR sketch rather than text and aim to obtain diverse 3D shapes that follow the input sketch structure.
@@ 3 Method
We present a conditional generation method that generates geometrically realistic shapes of a specific category conditioned on abstract, sparse, and inaccurate freehand VR sketches. Our goal is to enforce the generation to stay close to the input sketch (sketch fidelity) while providing sufficient diversity of 3D reconstructions.
@@ 3.1 Shape decoder
We represent 3D shapes using truncated signed distance functions (SDFs), as one of the most common 3D shape representations. This representation is limited to watertight meshes, but without loss of generality, here we assume that our meshes are watertight.
An SDF is a continuous function of the form:
@@
The auto-decoder is trained by minimizing an $L_{1}$ loss between the ground truth and predicted truncated signed distance values. The decoder
$\mathrm{D}_{\theta}([\boldsymbol{d}^{g},p_{i}])=\tilde{s}_{i},\ p_{i}\in\mathbb{R}^{3},\tilde{s}_{i}\in\mathbb{R}$ (2)
takes as input the 3D shape latent code $\boldsymbol{d}^{g}$ and the 3D point coordinates $p_{i}$; [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents a concatenation operation. The decoder predicts the per point signed distance value $\tilde{s}_{i}$. Once the decoder is trained, we freeze its parameters $\theta$.
@@
$\hat{\boldsymbol{{d}}^{g}}=\underset{\boldsymbol{d}^{g}}{\arg\min}\sum_{\left({p}_{i},{s}_{i}\right)\in G}\mathcal{L}\left(D_{\theta}\left(\boldsymbol{d}^{g},{p}_{i}\right),s_{i}\right)+\frac{1}{\sigma^{2}}\|\boldsymbol{d}^{g}\|_{2}^{2}$ (3)
where the latent vector $\boldsymbol{d}^{g}$ is initialized randomly, and the sum is taken over points of the 3D shape geometry $G$. We treat the estimated latent vector $\hat{\boldsymbol{{d}}^{g}}$ as a ground truth shape embedding and denote it as $\boldsymbol{d}^{g}$ for simplicity.
@@
Due to the sparsity of sketch inputs, we represent them as point clouds. We observe that if we only use sketches for training, we obtain very poor generalization to the test set. Therefore, we exploit a joint encoder for sketches and 3D shapes, using a PointNet++ [44] encoder. The encoder $\mathrm{E}_{\phi}(\cdot):\mathbb{R}^{N_{s}\times 3}\rightarrow\mathbb{R}^{256}$ embeds randomly sampled points from input shape surface $G\in\mathbb{R}^{N_{s}\times 3}$ or sketch strokes $C\in\mathbb{R}^{N_{s}\times 3}$ to a feature vector $\boldsymbol{e}^{g}\in\mathbb{R}^{256}$ and $\boldsymbol{e}^{c}\in\mathbb{R}^{256}$, respectively.
@@
First loss, $\mathcal{L}_{\text{shape}}$, operates only on 3D shape inputs, and ensures that the full network $\mathrm{D}_{\theta}([\mathrm{E}_{\phi}(\cdot),p_{i}])$ functions as a 3D shape autoencoder. First, we minimize the sum of $L_{1}$ losses between the truncated predicted and ground truth SDF values of sampled points $p_{i}\in\mathbb{R}^{3}$:
$\mathcal{L}_{\text{SDF}}(\phi)=\frac{1}{N_{s}}\sum_{p_{i}}|\text{tr}(\mathrm{D}_{\theta}([\boldsymbol{e}^{g},p_{i}]))-\text{tr}(s(p_{i}))|$ (5)
where the decoder parameters $\theta$ are from our auto-decoder and are frozen when training the sketch/shape encoder; $s(\cdot)$ denotes ground-truth 3D shape SDF values, and $\text{tr}(\cdot)\triangleq\min(\delta,\max(-\delta,\cdot))$.
Additionally, we ensure that the 3D shape embedding $\boldsymbol{e}^{g}=\mathrm{E}_{\phi}(f)$ maps directly to a latent representation of a 3D shape $\boldsymbol{d}^{g}$, computed with Eq. 3:
$\mathcal{L}_{g\text{-}L_{1}}(\phi)=|\boldsymbol{e}^{g}-\boldsymbol{d}^{g}|.$ (6)
We also found that adding contrastive latent term loss increases performance. Let’s assume that our mini-batch consists of $N_{g}$ shapes. First, we obtain latent representations $\{\boldsymbol{d}^{g}_{1},\dots,\boldsymbol{d}^{g}_{N_{g}}\}$ of 3D shapes, using Eq. 3. Then, we formulate our contrastive loss term as follows:
$\mathcal{L}_{\text{g-NCE}}(\phi)=-\sum_{i=1}^{N_{g}}\left[\log\frac{\exp\left(-\left|\boldsymbol{e}^{g}_{i}-\boldsymbol{d}^{g}_{i}\right|\right)}{\sum_{j=1}^{N_{g}}\exp\left(-\left|\boldsymbol{e}^{g}_{i}-\boldsymbol{d}^{g}_{j}\right|\right)}\right].$ (7)
@@
$\mathcal{L}_{\text{shape}}=\mathcal{L}_{\text{SDF}}(\phi)+\mathcal{L}_{g\text{-}L_{1}}(\phi)+\mathcal{L}_{g\text{-NCE}}(\phi).$ (8)
@@ 3.2.2 Sketch loss
Given the misalignment between sketches and reference 3D shapes in the dataset [35], as we show in Fig. 2, we aim for the reconstruction result to stay close to the sketch input. To achieve this goal, we design a sketch loss $\mathcal{L}_{\text{sketch}}$.
Since a sketch is a sparse representation of a 3D shape, the intended 3D shape surface should lie in the vicinity of sketch stroke points. Therefore, the reconstructed SDF values at those points should be close to zero. Formally, we define this loss as follows:
$\mathcal{L}_{\text{sketch}}(\phi)=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}|D(\boldsymbol{e}^{c},p^{c}_{i})|,$ (9)
where $p^{c}_{i}\in C$ is the $i$-th sample points from the conditioning sketch, and $N_{s}$ is the number of sampled points in a sketch.
@@ 3.2.3 Sketch-shape latent space alignment
The considered so far losses do not explicitly ensure that there is a meaningful mapping between sketches and 3D shapes’ latent representations. Therefore, we design additional losses encouraging an alignment in the feature space.
@@
$\mathcal{L}_{c\text{-NCE}}(\phi)=-\sum_{i=1}^{N_{c}}\left[\log\frac{\exp\left(-\left|\boldsymbol{e}^{c}_{i}-\boldsymbol{d}^{g}_{i}\right|\right)}{\sum_{j=1}^{N_{g}}\exp\left(-\left|\boldsymbol{e}^{c}_{i}-\boldsymbol{d}^{g}_{j}\right|\right)}\right],$ (10)
where $N_{c}$ is the number of sketches and $N_{g}$ is the number of shapes in the mini-batch. This loss pulls the encodings of a sketch and a reference shape closer than the encodings of a sketch and non-matching 3D shapes.
Additionally, we minimize the $L_{1}$ distance between the sketch $C$ embedding $\boldsymbol{e}^{c}=E_{\phi}(C)$ to the ground-truth shape latent code $\boldsymbol{d}^{g}$:
$\mathcal{L}_{c\text{-}L_{1}}(\phi)=|\boldsymbol{e}^{c}-\boldsymbol{d}^{g}|.$ (11)
@@
$\mathcal{L}_{\text{align}}=\mathcal{L}_{c\text{-}L_{1}}+\mathcal{L}_{c\text{-NCE}}$ (12)